OpenStudy (anonymous):

x1/3 * x1/4 can you please show work to show how this was done?

OpenStudy (anonymous):

Are those exponents

OpenStudy (anonymous):

No, it's just x and 1/3 times x and 1/4

OpenStudy (anonymous):

Is the following the problem expression?\[\left(\frac{1}{3}x\right)*\left(\frac{1}{4}x\right) \]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\left(\frac{1}{3}x\right)\left(\frac{1}{4}x\right)=\frac{x}{3}*\frac{x}{4}=\frac{x^2}{12} \]